Evaluate the following given that f(x) = sqrt of x+1 , g(x) = x^2-2 , and h(x) = (1)/(2x+3) . . solve (g+h)(-2) . . any takers?
Answer by fractalier(2101)

Okay,
(g+h)(-2) = ((x^2-2) + (1/(2x+3)))(-2) =
= (-2)^2 - 2 + (1/(2(-2)+3)) =
= 4 - 2 + 1/-1 =
= 4 - 2 - 1 =
= 1